Difference of proportions problem. Biofeedback is a treatment technique in which people are trained to improve their health by using signals from their own bodies. A study was conducted to compare a biofeedback treatment for chronic pain with an NSAID medical treatment. A group of 2100 newly diagnosed chronic pain patients were randomly assigned to receive one of the two treatments. After six weeks of treatments, the pain level of the patients was assessed and the results are shown in the table below:

Significant Reduction in Pain
TREATMENT Yes No Total
BioFeedback 580 450 1,030
NSAID 670 400 1,070
Total 1,240   840 2,100

d. Based on the result of a Null Hypothesis test that there is no difference in the proportion of patients answering YES between Biofeedback treatment and NSAID treatment, we would fail to reject the Null Hypothesis for a two-tailed test with an alpha level of .01.

Group of answer choices

True

False
Answer #1

Let the group which has answeres yes for biotreatment be 1 and the group which has answeres yes for NSAID be 2.

x1 = 580

x2 = 670

n1 = 1030

n2 = 1070

= 580/1030
= 0.5631

= 670/1070
= 0.6262

H0: p1 = p2

H1: p1 p2

Pooled propotion = (x1 + x2)/(n1 + n2)

= (580+670)/(1070+1030)

= 0.5952

Since x, n1-x1, x2 and n2-x2 are all > 5. We can approximate the difference between proportions by a normal distribution.

Hypothesis Testing:

p-value = 0.0032 < 0.01 i.e. H0 is rejected

Hence, the Correct answer is False.

Please upvote if you have liked my answer, would be of great help. Thank you.